Calculate the number of molecules in 0.020 mg Fe
1
Expert's answer
2020-10-23T09:16:21-0400

Molar mass of iron, M = 55.855 g/mol


0.020 mg of iron

= 0.020 × 10-3 g

= [0.020 × 10-3 g / 55.855 (g/mol)]

= 3.58 × 10-7 mol


In 1 mol of any substance number of molicule = 6.023×1023 mol-1


In 3.58 × 10-7 mol of substance number of molicule = 3.58 × 10-7 mol × 6.023×1023 mol-1

= 2.156 × 1017


Hence in 0.020 mg of Fe number of molicule = 2.156 × 1017